Assume that y varies directly as x. If y = -8 when x = -3, find x when y = 6. Leave your answer in fractional form
Answer by stanbon(75887)

Assume that y varies directly as x.
y = kx
----
Solve for "k" using "y = -8 when x = -3"
-8 = k*(-3)
k = 8/3
------
Equation:
y = (8/3)x
-------
Find x when y = 6.
---
6 = (8/3)*x
---
x = (3/8)6
x = 9/4
---------------
Cheers,
Stan H.